# Fast Training of Sinusoidal Neural Fields via Scaling Initialization

### v1

@@ 5.2 On selecting the optimal scaling factor
[…]
@@
Structural properties. In Fig. 6, we visualize how changing the data resolution ($512\times 512$ by default), model width ($512$ by default), and model depth ($4$ by default) affects the value of optimal scaling factor. We observe that the higher data resolution and shallower model depth lead to a need for a higher scaling factor, with the width having very weak correlation. Commonly, we find that the resolution and depth are both intimately related to the frequency spectrum; higher resolution leads to a need for a higher frequency, and deeper networks can effectively utilize the relatively smaller scaling factors to represent high-frequency signals.
@@ 6 Conclusion
In this paper, we revisit the traditional initialization scheme of sinusoidal neural fields, and propose a new method called weight scaling, which scales the initialized weights by a specific factor. In practice, WS significantly accelerates the training speed of SNF across various data modalities, even outperforming recent neural field training methods. We have discovered that the difference between the traditional and WS initialization comes from both the change in the frequency spectrum of the initial functional and the change in the layerwise learning rates, where the former plays a more crucial role in acceleration. Through extensive theoretical and empirical analyses, we have revealed that the WS positively impacts the learning dynamics by maintaining a well-conditioned optimization path throughout the training process.
Limitations and future direction. The limitations of our work lie in the fact that the analysis has been conducted for only neural fields using sinusoidal activation functions, rather than general neural field architectures; whether the weight scaling is effective on other architectures remains an open question. Also, from a theoretical perspective, our analyses does not provide a definitive answer on the acceleration mechanisms, lacking an explicit convergence bound. As a final remark, our work is closely related to the implicit bias of initialization scale (Azulay et al., 2021), which is an emerging topic in understanding the mechanics of neural network training. However, there remains a significant gap between theory and practice in this area, as the current interpretation is limited to shallow, homogeneous, and linear networks (Varre et al., 2023; Kunin et al., 2024), rather than general nonlinear networks. The next step is to understand the implicit bias in networks with general activation functions, where the scale of initialization is not necessarily equivalent to output scaling. We leave this as future work.
Reproducibility statement
Most experiments in our work are based on widely used neural field baselines, whose official source codes are publicly available online. The only new part, i.e., our proposed algorithm, is rather simple and only requires multiplying constants to the initial weights. While the implementation is quite simple, we will also release our code for our algorithm and the experiments as a public repository, upon the preparation of the camera-ready version of the paper.
@@
[…]
Appendix A Detailed statement and the derivation of 3.1
In this section, we provide a more detailed statement and derivation of 3.1, where we have extended the distribution-preserving claims of Sitzmann et al. (2020b) on $\alpha=1$ to the cases of $\alpha\geq 1$, thus covering the initialization with weight scaling.
We note that we generally follow the approach of Sitzmann et al. (2020b) in terms of the formalisms; we provide rough approximation guarantees (Lemmas A.1 and A.2), one of which relies on a numerical analysis of a function that is difficult to be expressed in a closed form (Lemma A.2).
@@
  • Activations to pre-activations. We show that, for $\alpha\geq 1$, the pre-activation of a layer with weight-scaled uniform activations and $\arcsin$ input distributions are approximately distributed as a Gaussian distribution with variance scaled by $\alpha$ (Lemma A.1).
  • Pre-activations to activations. Then, we show that, for $\alpha\geq 1$, post-activation distribution of the Gaussian distribution with $\alpha$-scaled variance is distributed as an $\arcsin$ distribution (Lemma A.2).
@@
Lemma A.1 (Arcsin to Gaussian distributions).
Let the weights $W_{l}$ of the $l$-th layer be sampled i.i.d. from $\mathrm{Unif}(-c/\sqrt{n},c/\sqrt{n})$, and the activation of the $(l-1)$-th layer $Z_{l-1}$ follow the i.i.d. $\mathrm{arcsin}(-1,1)$. Then, the distribution of the pre-activation of the $l$-th layer $Y_{l}$ converges in distribution to a Gaussian distribution $\mathcal{N}(0,c^{2}/6)$ as the number of hidden units $n$ tends to infinity.
Lemma A.2 (Gaussian to arcsin distributions, with numerical proof).
@@
Note that, by plugging in $c=\alpha\sqrt{6}$ to the Lemma A.1 as in the proposed weight scaling, we get the normal distribution $\mathcal{N}(0,\alpha^{2})$ as an output distribution.
@@
A.1 Proof of Lemma A.1
For simplicity, let us denote a pre-activation of a neuron in the $l$th layer by $y$. Let the weight vector connected to this output neuron as $\mathbf{w}\in\mathbb{R}^{n}$, with each entries drawn independently from $\mathrm{Unif}(-c/\sqrt{n},c/\sqrt{n})$. Also, let the corresponding input activation from the $(l-1)$th layer be denoted by $\mathbf{z}\in\mathbb{R}^{n}$, with each entry drawn independently from $\arcsin(-1,1)$. Now, it suffices to show that $y=\mathbf{w}^{\top}\mathbf{z}$ converges in distribution to $\mathcal{N}(0,c^{2}/6)$ as $n$ goes to infinity.
@@
Different from the settings of Sitzmann et al. (2020b), the sinusoidal activation is not a bijection in general within the support $[-3\alpha,3\alpha]$, i.e., not directly invertible. Thus, we break it down into three cases (Fig. 8): Case 1, where the endpoint $3\alpha$ is less than $\pi/2$, Case 2, where $3\alpha$ is in the increasing region of $\sin(\cdot)$, and Case 3, where $3\alpha$ is in the decreasing region. We study each case:
@@
where the last approximation invokes the Gaussian CDF approximation (Bowling et al., 2009). In this case, the CDF of $Y$ does not approximate the CDF of $\arcsin(-1,1)$. That is, the second-order derivative of $F$ decreases in $[-1,0]$ and increases in $[0,1]$, while the CDF of the $\arcsin$ distribution behaves in an opposite way.
Figure 8: Visualization of the cases in Lemma A.2. In $y=\sin(x)$, Red line : $x=3\alpha$ and Black line : $x=\frac{\pi}{2}$
Now, suppose that $\alpha$ is sufficiently large ($\alpha\geq\pi/6$), i.e., case 2 or case 3. Here, we will consider the nearest peak/valley that falls outside the $[-3\alpha,3\alpha]$ interval. More concretely, let $\gamma$ be defined as the period index of such peak/valley, i.e.,
@@
[…]
@@
Here, $F_{Y,\text{center}}(y)$ denotes the CDF of the bijective region, i.e., $x\in[-\pi/2,\pi/2]$, and $F_{Y,\text{inc}}(y)$ denotes the region where the $\sin(\cdot)$ is increasing and $F_{Y,\text{dec}}(y)$ denotes where $\sin(\cdot)$ is decreasing. Then, each of these components can be written as:
@@ Appendix B Proofs in Section 4
Lemma B.1 (From Yüce et al. (2022, Appendix A.2)).
@@
$f(x;\mathbf{W})=2\sum_{\ell\in\mathbb{Z}_{\text{odd}}}J_{\ell}(W^{(2)})\sin(\ell W^{(1)}x).$ (32)
B.2 Proof of Lemma 4.2
@@
First, we note that the first root of the $\ell$th order Bessel function, denoted by $j_{\ell,1}$, satisfies the inequality $j_{\ell,1}>\pi/2$ for all $\ell\in\mathbb{N}$. Next, recall the results of Ifantis & Siafarikas (1990) that
@@ C.1 Preliminaries
The Neural Tangent Kernel (NTK) framework views neural networks as kernel machines, where the kernel is defined as the dot-product of the gradients from two different data point (Jacot et al., 2018). In other words, each element of NTK can be expressed as $K(x_{i},x_{j})=\langle\nabla_{\mathbf{W}}f(x_{i}),\nabla_{\mathbf{W}}f(x_{j})\rangle$. Assuming an infinitely wide network trained with an infinitesimally small learning rate (i.e., gradient flow) and using a specific parameterization (i.e., NTK parameterization), the NTK converges to a deterministic kernel that does not evolve during training. However, in practice, neural networks are composed of finite-width layers, which introduces challenges for analysis using NTK-style theory. In this context, we study the empirical Neural Tangent Kernel (eNTK), where we do not require the assumption of a stationary kernel. Recent researches (Kopitkov & Indelman, 2020; Ortiz-Jiménez et al., 2021; Baratin et al., 2021) show that, with eNTKs non-linearly evolve during training, this kernel can serve as a measure of the difficulty of learning a particular task.
@@ C.2 Kernel regression with gradient flow
  • Another important aspect is understanding the characteristics of each direction. Empirical and theoretical evidence has shown that the NTK eigenfunctions corresponding to large eigenvalues are low-frequency functions. As a result, neural network training is biased towards reducing error in low-frequency regions (Ronen et al., 2019; Basri et al., 2020; Wang et al., 2021), providing strong evidence of spectral bias (Rahaman et al., 2019).
@@
C.3 Non-kernel approach: analysis via Hessian eigenvalues

### v3

@@ 5.2 On selecting the optimal scaling factor
$\underset{\alpha}{\text{maximize}}\quad$ $\mathrm{speed}(\alpha)\qquad\text{subject to}\quad\mathrm{TestPSNR}(\alpha)\geq 0.95\cdot\mathrm{TestPSNR}(1).$ (10)
@@
Structural properties. In Fig. 6, we visualize how changing the data resolution ($512\times 512$ by default), model width ($512$ by default), and model depth ($4$ by default) affects the value of optimal scaling factor. We observe that the higher data resolution and shallower model depth lead to a need for a higher scaling factor, with the width having very weak correlation. Commonly, we find that the resolution and depth are both intimately related to the frequency spectrum; higher resolution leads to a need for a higher frequency, and deeper networks can effectively utilize the relatively smaller scaling factors to represent high-frequency signals. In a similar context, we provide an analysis of the impact of the learning rate in Section E.3.
@@ 6 Conclusion
In this paper, we revisit the traditional initialization scheme of sinusoidal neural fields, and propose a new method called weight scaling (WS), which scales the initialized weights by a specific factor. In practice, WS significantly accelerates the training speed of SNF across various data modalities, even outperforming recent neural field methods. We have discovered that the difference between the traditional and WS initialization comes from both the change in the frequency spectrum of the initial functional and the change in the layerwise learning rates. Through extensive theoretical and empirical analyses, we have revealed that the WS positively impacts the learning dynamics by maintaining a well-conditioned optimization path throughout the training process.
Limitations and future directions. The limitations of our work lie in the fact that the analysis has been conducted for only neural fields using sinusoidal activation functions, rather than general neural field architectures; how the WS affects on other architectures remains an open question (see Section E.2). Also, from a theoretical perspective, our analyses does not provide a definitive answer on the acceleration mechanisms, lacking an explicit convergence bound. As a final remark, our work is closely related to the implicit bias of initialization scale (Azulay et al., 2021), which is an emerging topic in understanding the neural network training. However, there remains a significant gap between theory and practice in this area, as the current interpretation is limited to shallow and homogeneous (or linear) networks (Varre et al., 2023; Kunin et al., 2024). The next step is to understand the implicit bias in networks with general activation functions, where the scale of initialization is not necessarily equivalent to output scaling. Another research direction is to understand the effect of weight scaling in general neural field architectures (Section E.2), which can sometimes improve performance or have the opposite effect. We leave these for future work.
Acknowledgments
This work was supported in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00213710), in part by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (2022-0-00713), and in part by the NAVER-Intel Co-Lab; the work was conducted by POSTECH and reviewed by both NAVER and Intel.
@@ Acknowledgments
Appendix A Detailed statement and the derivation of 1
In this section, we provide a more detailed statement and derivation of 1, where we have extended the distribution-preserving claims of Sitzmann et al. (2020b) on $\alpha=1$ to the cases of $\alpha\geq 1$, thus covering the initialization with weight scaling.
We note that we generally follow the approach of Sitzmann et al. (2020b) in terms of the formalisms; we provide rough approximation guarantees (Lemmas 4 and 5), one of which relies on a numerical analysis of a function that is difficult to be expressed in a closed form (Lemma 5).
@@ Appendix A Detailed statement and the derivation of 1
  • Activations to pre-activations. We show that, for $\alpha\geq 1$, the pre-activation of a layer with weight-scaled uniform activations and $\arcsin$ input distributions are approximately distributed as a Gaussian distribution with variance scaled by $\alpha$ (Lemma 4).
  • Pre-activations to activations. Then, we show that, for $\alpha\geq 1$, post-activation distribution of the Gaussian distribution with $\alpha$-scaled variance is distributed as an $\arcsin$ distribution (Lemma 5).
@@
Lemma 4 (Arcsin to Gaussian distributions).
Let the weights $W_{l}$ of the $l$-th layer be sampled i.i.d. from $\mathrm{Unif}(-c/\sqrt{n},c/\sqrt{n})$, and the activation of the (l1)(l-1)-th layer $Z_{l-1}$ follow the i.i.d. $\mathrm{arcsin}(-1,1)$. Then, the distribution of the pre-activation of the $l$-th layer $Y_{l}$ converges in distribution to a Gaussian distribution $\mathcal{N}(0,c^{2}/6)$ as the number of hidden units $n$ tends to infinity.
Lemma 5 (Gaussian to arcsin distributions, with numerical proof).
@@
Note that, by plugging in $c=\alpha\sqrt{6}$ to the Lemma 4 as in the proposed weight scaling, we get the normal distribution $\mathcal{N}(0,\alpha^{2})$ as an output distribution.
@@
A.1 Proof of Lemma 4
For simplicity, let us denote a pre-activation of a neuron in the $l$th layer by $y$. Let the weight vector connected to this output neuron as $\mathbf{w}\in\mathbb{R}^{n}$, with each entries drawn independently from $\mathrm{Unif}(-c/\sqrt{n},c/\sqrt{n})$. Also, let the corresponding input activation from the (l1)(l-1)th layer be denoted by $\mathbf{z}\in\mathbb{R}^{n}$, with each entry drawn independently from $\arcsin(-1,1)$. Now, it suffices to show that $y=\mathbf{w}^{\top}\mathbf{z}$ converges in distribution to $\mathcal{N}(0,c^{2}/6)$ as $n$ goes to infinity.
@@ A.2 Proof of Lemma 5
Different from the settings of Sitzmann et al. (2020b), the sinusoidal activation is not a bijection in general within the support $3\alpha,3\alpha]$, i.e., not directly invertible. Thus, we break it down into three cases (Fig. 8): Case 1, where the endpoint $3\alpha$ is less than $\pi/2$, Case 2, where $3\alpha$ is in the increasing region of $\sin(\cdot)$, and Case 3, where $3\alpha$ is in the decreasing region. We study each case:
@@
where the last approximation invokes the Gaussian CDF approximation (Bowling et al., 2009). In this case, the CDF of $Y$ does not approximate the CDF of $\arcsin(-1,1)$. That is, the second-order derivative of $F$ decreases in [1,0][-1,0] and increases in $[0,1]$, while the CDF of the $\arcsin$ distribution behaves in an opposite way.
Figure 8: Visualization of the cases in Lemma 5. In $y=\sin(x)$, Red line : $x=3\alpha$ and Black line : $x=\frac{\pi}{2}$
Now, suppose that $\alpha$ is sufficiently large ($\alpha\geq\pi/6$), i.e., case 2 or case 3. Here, we will consider the nearest peak/valley that falls outside the $3\alpha,3\alpha]$ interval. More concretely, let $\gamma$ be defined as the period index of such peak/valley, i.e.,
@@
$\gamma_{1}$ $=\min\left\{n\in\mathbb{N}\mid\frac{(4n+1)\pi}{2}\geq 3\alpha~\cap~\sin\left(\frac{(4n+1)\pi}{2}\right)=+1\right\}\qquad{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{(peak; case 2)}}$ (16)
$\gamma_{2}$ $=\min\left\{n\in\mathbb{N}\mid\frac{(4n-1)\pi}{2}\geq 3\alpha~\cap~\sin\left(\frac{(4n-1)\pi}{2}\right)=-1\right\}\qquad{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{(valley; case 3)}}$ (17)
@@
Here, $F_{Y,\text{center}}(y)$ denotes the CDF of the bijective region, i.e., $x\in[-\pi/2,\pi/2]$, and $F_{Y,\text{inc}}(y)$ denotes the region where the $\sin(\cdot)$ is increasing and $F_{Y,\text{dec}}(y)$ denotes where $\sin(\cdot)$ is decreasing. Then, each of these components can be written as [^2] For notational brevity, we set $\alpha=1$ just for a moment. It will appear again in Eq. 23.:
@@ B.1 Proof of Lemma 2
Lemma 6 (From Yüce et al. (2022, Appendix A.2)).
@@
$f(x;\mathbf{W})=2W^{(3)}\sum_{\ell\in\mathbb{Z}_{\text{odd}}}J_{\ell}(W^{(2)})\sin(\ell W^{(1)}x).$ (32)
B.2 Proof of Lemma 3
@@
[…]
@@ C.1 Preliminaries
The Neural Tangent Kernel (NTK) framework views neural networks as kernel machines, where the kernel is defined as the dot-product of the gradients from two different data point (Jacot et al., 2018). In other words, each element of NTK can be expressed as $K(x_{i},x_{j})=\langle\nabla_{\mathbf{W}}f(x_{i}),\nabla_{\mathbf{W}}f(x_{j})\rangle$. Assuming an infinitely wide network trained with an infinitesimally small learning rate (i.e., gradient flow) and using a specific parameterization (i.e., NTK parameterization), the NTK converges to a deterministic kernel that does not evolve during training. However, in practice, neural networks are composed of finite-width layers, which introduces challenges for analysis using NTK-style theory. In this context, we study the empirical Neural Tangent Kernel (eNTK), where we do not require the assumption of a stationary kernel. Recent researches (Kopitkov and Indelman, 2020; Ortiz-Jiménez et al., 2021; Baratin et al., 2021) show that, with eNTKs non-linearly evolve during training, this kernel can serve as a measure of the difficulty of learning a particular task.
@@ C.2 Kernel regression with gradient flow
  • Another important aspect is understanding the characteristics of each direction. Empirical and theoretical evidence has shown that the NTK eigenfunctions corresponding to large eigenvalues are low-frequency functions. As a result, neural network training is biased towards reducing error in low-frequency regions (Basri et al., 2019; 2020; Wang et al., 2021), providing strong evidence of spectral bias (Rahaman et al., 2019).
@@
C.3 Insights into eNTK eigenvalue
In this subsection, we focus on the eNTK of two-layer sinusoidal networks with the scaling factor $\alpha$, i.e., $f(x;\mathbf{W})=W_{2}\sin(\alpha W_{1}x)$. The resulting kernel is a sum of two dot-product kernels
$K_{\text{NTK}}(x,x^{\prime})$ $=\underbrace{\langle\sin(\alpha W_{1}x),\sin(\alpha W_{1}x^{\prime})\rangle}_{K}+\underbrace{\langle\alpha W_{2}x\cos(\alpha W_{1}x),\alpha W_{2}x\cos(\alpha W_{1}x^{\prime})\rangle}_{K^{\prime}}.$ (52)
More specifically, we study the condition number and eigenvalue decay rate (EDR) of the kernel as follows.
  • Eigenvalue gap of eNTK. We start by approximating the sinusoidal eNTK as a Gaussian kernel (Rahimi and Recht, 2007; Mehta et al., 2021), which allows for an analytic form of the spectral decomposition. Using this approximation, we can also estimate the EDR. Here, we observe that an increase in the scaling factor $\alpha$ directly leads to a decrease in the EDR.
  • Condition number. Building on recent studies of the condition number of the NTK, which significantly affects the convergence rate of neural networks (particularly in the kernel regime) (Xiao et al., 2020; Liu and Hui, 2023), we demonstrate that WS-SNF approximately achieves a low condition number for the eNTK, with the scaling factor $\alpha$ playing a crucial role due to the lower EDR.
We start with the results from Mehta et al. (2021).
Theorem 7 (Theorem 1 in Mehta et al. (2021).).
Suppose each entry of $W_{1}$ is initialized with a Gaussian distribution of mean 0 and variance $\sigma^{2}$. Then for any $x$ and $x^{\prime}$, we have
$\left|\mathbb{E}_{W_{1}}\left[\langle\sin(W_{1}x),\sin(W_{1}x^{\prime})\rangle\right]\right|\leq C_{1}\exp\left(-\frac{\sigma^{2}\|x-x^{\prime}\|^{2}_{2}}{2}\right),$ (53)
Theorem 7 states that the expectation over the first-layer weights can be bounded by Gaussian kernels, where the bandwidth is determined by the initialization variance $\sigma$. The scaling factor in Eq. 52 directly affects the variance term, which can be replaced with $\alpha^{2}\sigma^{2}$, without loss of generality. Note that it can be handled similarly in the case of cosine activation function.
Returning to the eNTK, we start by analyzing $K$. From Eq. 53, the kernel element can be approximated as
$K(x_{i},x_{j})\leq C_{1}\exp\left(-\frac{\alpha^{2}\sigma^{2}\|x_{i}-x_{j}\|^{2}}{2}\right)=K^{*}(x_{i},x_{j}).$ (54)
Assuming $K^{*}-K\succeq 0$, then $\lambda_{i}^{*}\geq\lambda_{i}$ holds for all eigenvalue indices $i$, where $\lambda_{i}^{*}$ and $\lambda_{i}$ denote the $i$-th eigenvalues of the kernels $K^{*}$ and $K$, respectively. Next step is to compute the eigenvalues of $K^{*}$ (i.e., eigenvalue upper bound of the $K$). The eigenvalues of $K^{*}$ can be represented as (Rasmussen and Williams, 2006):
$\lambda_{i}\leq\lambda_{i}^{*}=C_{1}\sqrt{\frac{2}{1+2\alpha^{2}\sigma^{2}+\sqrt{1+4\alpha^{2}\sigma^{2}}}}\left(\frac{2\alpha^{2}\sigma^{2}}{1+2\alpha^{2}\sigma^{2}+\sqrt{1+4\alpha^{2}\sigma^{2}}}\right)^{i},$ (55)
and $K^{\prime}$ can be handled similarly. In Eq. 55, we assume that $x\sim\mathcal{N}(0,\sigma_{x}^{2})$, and set $\sigma_{x}=1$ without loss of generality. Note that the eigenfunctions of the Gaussian kernel $K^{*}$ can be computed via Hermite polynomial expansion; however, this does not guarantee characteristics of the eigenfunctions of $K$, so we do not take into account. In Eq. 55, we can compute the eigenvalue decay rate (EDR) from arbitrary index $k$ to $k+1$ by examining the base number. Intuitively, we can see that the larger the $\alpha$, the smaller the EDR. Considering both kernels $K$ and $K^{\prime}$, characterizing the exact eigenvalue of $K_{\text{NTK}}=K+K^{\prime}$ is open problem in the field of random matrix theory (Tao, 2012). However, we analyze the spectrum indirectly through Weyl’s inequality from Courant-Fischer’s theorem (Horn and Johnson, 2012). Let $\lambda_{k}(K)$ and $\lambda_{k}(K^{\prime})$ the $k(\leq n)$-th largest eigenvalue of each kernel. Then:
$\lambda_{k}(K^{\prime})+\lambda_{1}(K)\rho^{n-1}\leq\lambda_{k}(K_{\text{NTK}})\leq\lambda_{k}(K^{\prime})+\lambda_{1}(K).$ (56)
Here, $\rho$ denotes the EDR of $K$. In other words, the bound of $\lambda_{k}(K_{\text{NTK}})$ is determined by the EDR of $K$, where a lower EDR (i.e., a larger scaling factor) contributes to an increase in the lower bound of the eNTK eigenvalue for each index.
This theoretical analysis aligns with our empirical observations. Specifically, as shown in Fig. 5(b), the $x$-axis represents the distribution of normalized eigenvalues. The eigenvalues of the weight-scaled network are more densely distributed compared to the default case (i.e., no weight scaling applied), indicating that the EDR of the weight-scaled network is lower. Intuitively, a smaller EDR corresponds to a smaller condition number, further suggesting that the optimization trajectory of weight scaling is relatively well-conditioned.
Recently, Saratchandran et al. (2024a) theoretically derived the minimum eigenvalue of the eNTK in periodically activated MLPs of general depth. Under certain assumptions, the minimum eigenvalue is shown to be asymptotically bounded, involving the variance of the initialization distribution (specifically, larger the variance, larger the minimum eigenvalue), which also supports our results.
C.4 Non-kernel approach: analysis via Hessian eigenvalues